# Cesàro-Hardy operators on $L^{p}[0,1]$: fine spectrum, weighted Koopman semigroups and invariant subspaces

### 22 Apr 2026 v1

@@ 1. Introduction
The discrete Cesàro-Hardy operator’s spectrum and fine spectra have similarly been analyzed in various spaces, including $\ell_{p}$ $(1<p<\infty)$ [26, 19], $c_{0}$ [38, 4], $\ell_{\infty}$, Bachelis spaces $N_{p}$ $(1<p<\infty)$ [21], $bv_{0}$, and bvbv [35, 36], as well as in weighted $\ell_{p}$ spaces [10, 7], discrete Cesàro spaces $\operatorname{ces}_{p}$ $(p\in\{0\}\cup(1,\infty))$ [20], and their dual spaces $d_{s}$ $(1<s<\infty)$ [14].
@@
where $\chi_{I}(s)$ denotes the characteristic function of the interval II. The operator families, $(T(t))_{t\geq 0}$ and $(S(t))_{t\geq 0}$ given by
@@
are $C_{0}$-semigroups on $L^{p}[0,1]$ for $1\leq p<\infty$, see Theorem $2.$ Both semigroups are referred to in the literature as Koopman or composition semigroups. Next, with this representation in mind, we obtain the fine spectrum of both infinitesimal generators and transfer their spectral properties, first to the semigroup, using standard $C_{0}$-semigroup theory, see Theorems 4.1 and 4.2. And later on, to the Cesàro-Hardy operator using the Hille-Phillips functional calculus and the functional calculus for sectorial operators developed by M. Haase in [27], see Theorems 5.3 and 5.4. This semigroup approach allows us to get, in a unified way, the spectrum of wide class of integral operators, see Section 5.
@@
While we initially presented the work of [31] as the motivation for addressing the problem discussed in this paper, it is important to acknowledge that the earliest known reference establishing the connection between $C_{0}$-semigroups and the Cesàro-Hardy operator is the article by $C.$ Cowen [18]. Afterward, the technique was further employed by A. Arvanitidis and A. Siskakis [13], who proved that the half-plane versions of Cesàro-Hardy operators on Hardy spaces are bounded. Similarly, W. Arendt and B. de Pagter [12] examined the Cesàro-Hardy operator defined in an interpolation space $E$ between $L^{1}$ and $L^{\infty}$ on the positive real half-line. Additionally, as we mentioned before, the generalized discrete Cesàro-Hardy operator was studied in [1], and more recently in [2] the authors used the $C_{0}$-semigroup approach to determine the fine spectrum of generalized Hausdorff operators on various analytic spaces, including a generalized Cesàro-Hardy operator.
@@
To conclude this introduction, we recall some classical definitions. Let $X$ be a Banach space, $\mathcal{B}(X)$ will denote the Banach algebra of bounded linear operators on $X$ with the operator norm $\left\|\cdot\right\|$. Let $1\leq p\leq\infty$, the Lebesgue spaces $L^{p}[0,1]$ are Banach spaces with the $L^{p}$-norm denoted by $\left\|\cdot\right\|_{p}$. For $1\leq p<\infty$ the dual space of $L^{p}$ is $L^{p^{\prime}}$, where $p^{\prime}$ denotes the conjugate exponent of pp. For $1<p<\infty$, $p^{\prime}$ satisfies $\frac{1}{p}+\frac{1}{p^{\prime}}=1$. If $p=1$, then $p^{\prime}=\infty$ and $\frac{1}{p^{\prime}}$ is understood as 0. Throughout the paper we will maintain this notation.
@@ 2. Weighted Koopman semigroups on $L^{p}[0,1]$
Theorem $2.$
Let $1\leq p<\infty$. The one-parameter operator families $(T(t))_{t\geq 0}$ and $(S(t))_{t\geq 0}$ are $C_{0}$-semigroups on $L^{p}[0,1]$ with growth bounds $1/p^{\prime}$ and 1/p-1/p, respectively. Their infinitesimal generators, AA and $B$, are given by
@@ 3. Spectra of differential operators
In this section we present the fine spectrum of the infinitesimal generators AA and $B$ on $L^{p}[0,1].$
@@
Let $1\leq p<\infty.$ The point spectrum of $B$ is empty, that is, $\sigma_{point}(B)=\emptyset.$ The point spectrum of AA is the set $\sigma_{point}(A)=\{\lambda\in\mathbb{C}\,:\,{\mathfrak{R}}{\mathfrak{e}}\,\lambda<-1/p^{\prime}\}.$ Furthermore, the eigenspace of each $\lambda\in\sigma_{point}(A)$ is one-dimensional and generated by $g_{\lambda}(s):=\frac{1}{s^{\lambda+1}}.$
@@
Let $1\leq p<\infty.$ The spectrum of AA is the set $\sigma(A)=\{\lambda\in\mathbb{C}\,:\,{\mathfrak{R}}{\mathfrak{e}}\,\lambda\leq-1/p^{\prime}\}.$ Moreover, for $\lambda\in\mathbb{C}$ with ${\mathfrak{R}}{\mathfrak{e}}\,\lambda>-1/p^{\prime},$
@@
Let ${\mathfrak{R}}{\mathfrak{e}}\,\lambda>-1/p^{\prime}$, by the integral representation of the resolvent operator of AA in terms of the semigroup $(T(t))_{t\geq 0}$ (see e.g. [23, Th. II.1.10]) and Proposition 2.5, one has
@@
Let $1\leq p<\infty.$ The differential operator AA satisfies that
@@ 4. Spectra of Koopman semigroups
At Section 2 we introduced the operator families $(T(t))_{t\geq 0}$ and $(S(t))_{t\geq 0}$ which are $C_{0}$-semigroups for $1\leq p<\infty,$ and which have the operators AA and $B$ as infinitesimal generators respectively. In this section, we use the spectral study of the generators in Section 3 in order to study the spectral properties of the $C_{0}$-semigroups $(T(t))_{t\geq 0}$ and $(S(t))_{t\geq 0}$ defined on $L^{p}[0,1]$ for $1\leq p<\infty$. From now on, we denote by $B(c,r)$ the open ball with center $c$ and radius $r$. Also, from now on, the following subspaces
@@ 7. Universality and closed invariant subspaces for the semigroup $T(t)$
Remark $7.$
It is known that for a contraction $C_{0}$-semigroup, the lattice of the closed invariant subspaces of the semigroup is the same that the lattice of the closed invariant subspaces of the cogenerator (see [25, Chap. 2, Theorem 10.9]). In particular, $(e^{t}T(2t))_{t\geq 0}$ is contraction $C_{0}$-semigroup on $L^{2}[0,1]$ generated by $2A+I,$ and whose cogenerator is $V=((2A+I)+I)((2A+I)-I)^{-1},$ see Proposition 2.1 and Theorem $2.$ Therefore, by Remark 5.5 one can write

### 23 Apr 2026 v2

@@ 1. Introduction
The discrete Cesàro-Hardy operator’s spectrum and fine spectra have similarly been analyzed in various spaces, including $\ell_{p}$ $(1<p<\infty)$ [26, 19], $c_{0}$ [38, 4], $\ell_{\infty}$, Bachelis spaces $N_{p}$ $(1<p<\infty)$ [21], $bv_{0}$, and $bv$ [35, 36], as well as in weighted $\ell_{p}$ spaces [10, 7], discrete Cesàro spaces $\operatorname{ces}_{p}$ $(p\in\{0\}\cup(1,\infty))$ [20], and their dual spaces $d_{s}$ $(1<s<\infty)$ [14].
@@
where $\chi_{I}(s)$ denotes the characteristic function of the interval $I$. The operator families, $(T(t))_{t\geq 0}$ and $(S(t))_{t\geq 0}$ given by
@@
are $C_{0}$-semigroups on $L^{p}[0,1]$ for $1\leq p<\infty$, see Theorem 2.2. Both semigroups are referred to in the literature as Koopman or composition semigroups. Next, with this representation in mind, we obtain the fine spectrum of both infinitesimal generators and transfer their spectral properties, first to the semigroup, using standard $C_{0}$-semigroup theory, see Theorems 4.1 and 4.2. And later on, to the Cesàro-Hardy operator using the Hille-Phillips functional calculus and the functional calculus for sectorial operators developed by M. Haase in [27], see Theorems 5.3 and 5.4. This semigroup approach allows us to get, in a unified way, the spectrum of wide class of integral operators, see Section 5.
@@
While we initially presented the work of [31] as the motivation for addressing the problem discussed in this paper, it is important to acknowledge that the earliest known reference establishing the connection between $C_{0}$-semigroups and the Cesàro-Hardy operator is the article by C.C. Cowen [18]. Afterward, the technique was further employed by A. Arvanitidis and A. Siskakis [13], who proved that the half-plane versions of Cesàro-Hardy operators on Hardy spaces are bounded. Similarly, W. Arendt and B. de Pagter [12] examined the Cesàro-Hardy operator defined in an interpolation space $E$ between $L^{1}$ and $L^{\infty}$ on the positive real half-line. Additionally, as we mentioned before, the generalized discrete Cesàro-Hardy operator was studied in [1], and more recently in [2] the authors used the $C_{0}$-semigroup approach to determine the fine spectrum of generalized Hausdorff operators on various analytic spaces, including a generalized Cesàro-Hardy operator.
@@
To conclude this introduction, we recall some classical definitions. Let $X$ be a Banach space, $\mathcal{B}(X)$ will denote the Banach algebra of bounded linear operators on $X$ with the operator norm $\left\|\cdot\right\|$. Let $1\leq p\leq\infty$, the Lebesgue spaces $L^{p}[0,1]$ are Banach spaces with the $L^{p}$-norm denoted by $\left\|\cdot\right\|_{p}$. For $1\leq p<\infty$ the dual space of $L^{p}$ is $L^{p^{\prime}}$, where $p^{\prime}$ denotes the conjugate exponent of $p$. For $1<p<\infty$, $p^{\prime}$ satisfies $\frac{1}{p}+\frac{1}{p^{\prime}}=1$. If $p=1$, then $p^{\prime}=\infty$ and $\frac{1}{p^{\prime}}$ is understood as 0. Throughout the paper we will maintain this notation.
@@ 2. Weighted Koopman semigroups on $L^{p}[0,1]$
Theorem 2.2.
Let $1\leq p<\infty$. The one-parameter operator families $(T(t))_{t\geq 0}$ and $(S(t))_{t\geq 0}$ are $C_{0}$-semigroups on $L^{p}[0,1]$ with growth bounds $1/p^{\prime}$ and 1/p-1/p, respectively. Their infinitesimal generators, $A$ and $B$, are given by
@@ 3. Spectra of differential operators
In this section we present the fine spectrum of the infinitesimal generators $A$ and $B$ on $L^{p}[0,1].$
@@
Let $1\leq p<\infty.$ The point spectrum of $B$ is empty, that is, $\sigma_{point}(B)=\emptyset.$ The point spectrum of $A$ is the set $\sigma_{point}(A)=\{\lambda\in\mathbb{C}\,:\,{\mathfrak{R}}{\mathfrak{e}}\,\lambda<-1/p^{\prime}\}.$ Furthermore, the eigenspace of each $\lambda\in\sigma_{point}(A)$ is one-dimensional and generated by $g_{\lambda}(s):=\frac{1}{s^{\lambda+1}}.$
@@
Let $1\leq p<\infty.$ The spectrum of $A$ is the set $\sigma(A)=\{\lambda\in\mathbb{C}\,:\,{\mathfrak{R}}{\mathfrak{e}}\,\lambda\leq-1/p^{\prime}\}.$ Moreover, for $\lambda\in\mathbb{C}$ with ${\mathfrak{R}}{\mathfrak{e}}\,\lambda>-1/p^{\prime},$
@@
Let ${\mathfrak{R}}{\mathfrak{e}}\,\lambda>-1/p^{\prime}$, by the integral representation of the resolvent operator of $A$ in terms of the semigroup $(T(t))_{t\geq 0}$ (see e.g. [23, Th. II.1.10]) and Proposition 2.5, one has
@@
Let $1\leq p<\infty.$ The differential operator $A$ satisfies that
@@ 4. Spectra of Koopman semigroups
At Section 2 we introduced the operator families $(T(t))_{t\geq 0}$ and $(S(t))_{t\geq 0}$ which are $C_{0}$-semigroups for $1\leq p<\infty,$ and which have the operators $A$ and $B$ as infinitesimal generators respectively. In this section, we use the spectral study of the generators in Section 3 in order to study the spectral properties of the $C_{0}$-semigroups $(T(t))_{t\geq 0}$ and $(S(t))_{t\geq 0}$ defined on $L^{p}[0,1]$ for $1\leq p<\infty$. From now on, we denote by $B(c,r)$ the open ball with center $c$ and radius $r$. Also, from now on, the following subspaces
@@ 7. Universality and closed invariant subspaces for the semigroup $T(t)$
Remark 7.7.
It is known that for a contraction $C_{0}$-semigroup, the lattice of the closed invariant subspaces of the semigroup is the same that the lattice of the closed invariant subspaces of the cogenerator (see [25, Chap. 2, Theorem 10.9]). In particular, $(e^{t}T(2t))_{t\geq 0}$ is contraction $C_{0}$-semigroup on $L^{2}[0,1]$ generated by $2A+I,$ and whose cogenerator is $V=((2A+I)+I)((2A+I)-I)^{-1},$ see Proposition 2.1 and Theorem 2.2. Therefore, by Remark 5.5 one can write